OpenStudy (maheshmeghwal9):

\[\frac{\partial^2 \psi}{\partial x^2}+\frac{\partial^2 \psi}{\partial y^2}+\frac{\partial^2 \psi}{\partial z^2}+\frac{8 \pi ^2 m}{h^2}(E-V) \psi =0.\] Shrodinger Equation.

OpenStudy (maheshmeghwal9):

wt does this actually mean? wt are the solutions?

OpenStudy (unklerhaukus):

wave function are solutions

OpenStudy (maheshmeghwal9):

so \[\psi \space; \space \psi ^2\] are solutions ??????

OpenStudy (unklerhaukus):

do you want a text to read?

OpenStudy (maheshmeghwal9):

no; i have read already but couldn't understand. but wanna understand from helpers as u:)

OpenStudy (unklerhaukus):

The time dependent wavefunction \[\Psi(x,t)=\psi(x)\phi(t)\] is normalized \[\int\limits_{-\infty}^\infty\left|\Psi(x,t)\right|^2\text dx=1\]

OpenStudy (maheshmeghwal9):

wt is the meaning of this ?

OpenStudy (maheshmeghwal9):

i think u have broken the function as a product of 2 other functions but wt about the integral?

OpenStudy (unklerhaukus):

you have quoted thee dimensional time independent schrödinger equation

OpenStudy (unklerhaukus):

the mod square of Psi is Psi times its complex conjugate, \[\left|\Psi\right|^2=\Psi^*\Psi\] the time variable will go because \[\psi(t)=e^{iEt/\hbar}\]

OpenStudy (unklerhaukus):

\[\sum\limits_{j=0}^\infty P_j=1\]the sum of probabilities is equal to one \[\int\limits_{-\infty}^{\infty}|\Psi(x,t)|^2=\int\limits_{-\infty}^{\infty}\rho(x)=1\]

OpenStudy (maheshmeghwal9):

so this equation contains infinite solution but when broken in two parts have wave function as its solution?????????? Am i right????????

OpenStudy (unklerhaukus):

i think you are right , the psi's are the solutions yes

OpenStudy (anonymous):

very few people actually know whta he hel it is rest are just memorizing and using it and i hate it. . . :P because me too can't understand it because no one is really there to explain n simple words. . . i would advice you to put this question in physics forum.

OpenStudy (maheshmeghwal9):

ok thanx a lot but plz tell me how this equation works in finding orbitals?

OpenStudy (unklerhaukus):

orbitals are the shapes of the solution

OpenStudy (maheshmeghwal9):

oh i see thanx once again:)

OpenStudy (anonymous):

this equation expresses potential energy of the patical as well kinetic energy wrt all theree dimensions that is x, y, and z. . . this equation contains all these terms. . .

OpenStudy (unklerhaukus):

of course there are infinite solutions but there is higher probability density where the cloud of solutions is less faint(dense)

OpenStudy (unklerhaukus):

less faint ; more dense

OpenStudy (maheshmeghwal9):

ok i gt so \[\psi^2 \] is our probility density:)

OpenStudy (unklerhaukus):

\[\rho(x,y,z) \]is this spacial density

OpenStudy (unklerhaukus):

your pretty close it is the mod square of Psi \[|\Psi(t,x,y,x)|=\Psi(t,x,y,x)^*\Psi(t,x,y,x)\]

OpenStudy (unklerhaukus):

=\(\rho(x,y,z)\)

OpenStudy (maheshmeghwal9):

oh ok i gt thanx:)

OpenStudy (unklerhaukus):

QM uses \(a^*\) to denote the complex conjugate of \(a\) you might be more familiar with \(\overline a\)

OpenStudy (unklerhaukus):

if \(z=a+ib\)\[\overline z=z^*=a-ib\]

OpenStudy (maheshmeghwal9):

ok!